OpenStudy (anonymous):

write the equation of the line that is parallel to y=-x+4 and goes through the points (-3,-2)

OpenStudy (nincompoop):

identify the slope of the line you are given then use that slope to obtain the equation of the other line with the given points

OpenStudy (nincompoop):

do you know the slope of the original line?

OpenStudy (anonymous):

the equation is y=-x=4... would the slope 4 over 1??

OpenStudy (anonymous):

(geometry)

OpenStudy (nincompoop):

where did you get that? y = mx+b in that slope-intercept form the m in "mx" is your slope, and the y-intercept (0,y) is the b

OpenStudy (anonymous):

the information it gives me, write the equation of the line that is parallel to y=--x=4 and goes through the points (-3,-2)

OpenStudy (nincompoop):

this was the equation of your first line y=-x+4 that means your slope is -1 and that your y-intercept is (0,4) now to look for a line parallel to that with given points (-3, -2), you need to use the slope of -1 in the point-slope form y-y1=m(x-x1) it becomes y-(-2)=-1[x-(-3)]

OpenStudy (nincompoop):

just solve, simplify and you should be done if you are being asked to put it in the slope-intercept form, then do it.

OpenStudy (nincompoop):

@punyeta has something profound to add .. I can't wait to find out what it is

OpenStudy (anonymous):

x+y= -5

OpenStudy (anonymous):

so the answer would be y-(-2)=-1[x-(-3)] for the problem?

OpenStudy (nincompoop):

that answer is not yet simplified y-(-2) amounts to y+2

OpenStudy (nincompoop):

use distributive property with -1[x-(-3)]

OpenStudy (nincompoop):

is this your first time taking algebra?

OpenStudy (anonymous):

taking geometry

OpenStudy (nincompoop):

or geometry... to me they're fundamentally integrated and cannot be separated

OpenStudy (anonymous):

y-(-2)=-1[x-(-3)] y+2= -1(x+3) y+2=-x-3 x+y=-3-2 x+y=-5

OpenStudy (nincompoop):

ax+by=c is the standard form of a line y=mx+b or some other countries use y=mx+c for the y-intercept form.

OpenStudy (anonymous):

ok

OpenStudy (nincompoop):

I don't think you understood anything, @wildpuppy

OpenStudy (nincompoop):

post another problem and I want to see you solve it yourself

OpenStudy (anonymous):

that was the only question that was complicated like that one

OpenStudy (nincompoop):

that wasn't complicated at all learn what x-intercept and y-intercepts are know what a slope is and different forms of equation of a line

OpenStudy (anonymous):



OpenStudy (nincompoop):

close this thread and then write a new question?

OpenStudy (anonymous):

i did, i posted it

OpenStudy (nincompoop):

where? tag me

OpenStudy (anonymous):

heres the link http://openstudy.com/study#/updates/526d6976e4b0ff172476619b

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.